OpenStudy (anonymous):

PLEASE HELP. PLEASE. Medal and FAN!!!! I need to figure this out. The radioisotope radon-222 has half-life of 3.8 days. how much of a 73.9-g sample of radon-222 would be left after approximately 23 days?

OpenStudy (quickthing):

Let me see..

OpenStudy (quickthing):

A(f) = A(o) x 1 /2^n But here were already given the half life. So the question becomes how many half lives occur over 23 days? (23/3.8) is about 6 A(f) = 73.9 x 1/2^6 A(f) = 73.9 * 1 / 64 or about 1.15 grams

OpenStudy (quickthing):

There you go

OpenStudy (anonymous):

OMG. Thank you so much!!!!!